# Designing Phase Sensitive Probes of Monopole Superconducting Order

### v1

@@ E.1 $q_{p}=-1$ monopole superconductor
Figure 9: Magnitude and phase winding of the $q_{p}=1$ monopole superconducting gap function, projected onto the Fermi surface of the tight-binding model. Presented is the lattice analogue to $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}(\tilde{\mathbf{k}})$ in Eq. (24). (a-b) The magnitude of the gap function. (c-d) Phase winding of the pairing order, chosen in a gauge which is well-defined at the north or south pole respectively. Results are computed using the same parameters as Fig. 8.
E.2 $q_{p}=3$ monopole superconductor
@@
We similarly construct a tight binding model for the monopole superconducting order having pair monopole charge $q_{p}=3$, as introduced in Sec. IV.3. Here, the band Hamiltonian describes a magnetic doped triple Weyl semimetal. As $C_{6}$ symmetry is needed to stabilize the triple Weyl nodes, the tight-binding model is defined on a three-dimensional hexagonal lattice with $D_{6h}$ symmetry, which consists of $AA$ stacked two-dimensional triangular lattices.
@@
$H_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}=\sum_{\mathbf{k}}\sum_{\sigma,\sigma^{\prime}=\uparrow,\downarrow}c^{\dagger}_{\mathbf{k},\sigma}\left[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}(\mathbf{k})\right]_{\sigma,\sigma^{\prime}}c_{\mathbf{k},\sigma^{\prime}}+\sum_{\mathbf{k}}\sum_{\sigma,\sigma^{\prime}=\uparrow,\downarrow}\Big{(}c^{\dagger}_{\mathbf{k},\sigma}\big{[}\Delta_{\mathrm{inter}}(\mathbf{k})\big{]}_{\sigma,\sigma^{\prime}}c^{\dagger}_{-\mathbf{k},\sigma^{\prime}}+\mathrm{h.c.}\Big{)},$ (83)
with $\Delta_{\mathrm{inter}}(\mathbf{k})$ being the inter-Fermi surface pairing. Above, the Fourier transform of the band Hamiltonian kernel for the triple Weyl semimetal is
@@
in which $g_{x}(\mathbf{k})=2t_{1}(\sin k_{1}-\sin k_{2}+\sin k_{3})$, $g_{y}(\mathbf{k})=({-2t_{2}}/{3\sqrt{3}})[\sin(k_{1}+k_{2})-\sin(k_{2}+k_{3})+\sin(k_{3}-k_{1})]$, and $g_{z}(\mathbf{k})={2t_{z}(\cos K_{0,z}-\cos k_{z})}+{2m(3-\cos k_{1}-\cos k_{2}-\cos k_{3})}$. The above system hosts two triple Weyl nodes pinned along the $k_{z}$ axis at ±$\pm\mathbf{K}_{0}=\left(0,0,\pm K_{0,z}\right)$ with opposite winding numbers $\mp 3$. In the low-energy regime, the above tight-binding model of the band Hamiltonian reduces to that of the continuum model given in Eq. (36). Nearest and next-nearest neighbor hopping amplitudes are denoted by $t_{1}$ and $t_{2}$ respectively which, for simplicity, we choose to be $t_{1}=t_{2}=8t_{z}\sin K_{0,z}$ to compare to the continuum model. Above, $k_{i}\equiv\bm{\delta}_{i}\cdot\mathbf{k}$, where $\bm{\delta}_{1}=(1,0,0)$, $\bm{\delta}_{2}=(\frac{1}{2},\frac{\sqrt{3}}{2},0)$, and $\bm{\delta}_{3}=(-\frac{1}{2},\frac{\sqrt{3}}{2},0)$ are the lattice vectors connecting neighbouring sites in the $ab$-plane, and the lattice constant has been set to unity. At appropriate doping, the above system realizes two disjoint Fermi surfaces, $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$, situated about triple Weyl nodes at ±$\pm\mathbf{K}_{0}$ and related by parity, $\mathcal{H}_{\mathrm{kin},1}^{(\nu=-3)}(\mathbf{k})=\sigma_{z}\mathcal{H}_{\mathrm{kin},2}^{(\nu=+3)}(-\mathbf{k})\sigma_{z}$.
@@
The nonzero matrix elements, which feature nearest and next-nearest neighbour hopping, read
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}}=(2t_{z}\cos K_{0,z}+6m)\sigma_{z}-\mu\sigma_{0},$ (85a)
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{1}}=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{1};\mathbf{n}}^{\dagger}=-m\sigma_{z}-it_{1}\sigma_{x}$ (85b)
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{2}}=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{2};\mathbf{n}}^{\dagger}=-m\sigma_{z}+it_{1}\sigma_{x}$ (85c)
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{3}}=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{3};\mathbf{n}}^{\dagger}=-m\sigma_{z}-it_{1}\sigma_{x}$ (85d)
@@
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{1}+\bm{\delta}_{2}}=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{1}+\bm{\delta}_{2};\mathbf{n}}^{\dagger}=-\frac{it_{2}\nu}{3\sqrt{3}}\sigma_{y}$ (85e)
@@
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{2}+\bm{\delta}_{3}}=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{2}+\bm{\delta}_{3};\mathbf{n}}^{\dagger}=\frac{it_{2}\nu}{3\sqrt{3}}\sigma_{y}$ (85f)
@@
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{3}-\bm{\delta}_{1}}=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{3}-\bm{\delta}_{1};\mathbf{n}}^{\dagger}=-\frac{it_{2}\nu}{3\sqrt{3}}\sigma_{y}$ (85g)
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{z}}=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{WSM}}]_{\mathbf{n}+\bm{\delta}_{z};\mathbf{n}}^{\dagger}=-t_{z}\sigma_{z}.$ (85h)
with $\bm{\delta}_{i=1,2,3}$ and $\mathbf{\delta}_{z}=(0,0,1)$ being the lattice vectors connecting nearest neighbours. When there is inter-Fermi surface $s$-wave pairing, as in Eq. (82), the system gives rise to the monopole superconducting order $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ confined to the topological sector $q_{p}=3$ and with global angular momentum $l_{z,\mathrm{glob}}=0$, as described in Sec. IV.3 in the main text.
Figure 10: Magnitude and phase winding of the $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ monopole superconducting pairing order computed with and projected onto the band Fermi surface of the tight-binding model in Eq. (84). (a-b) The magnitude of the gap function, exhibiting point nodes at the north and south pole. (c-d) Phase winding of the pairing order, chosen in a gauge with a Dirac string located at the south and north poles respectively. Parameters are chosen as $\mu=0.1$, $K_{0}=\pi/3$, $t_{xy}=t_{z}\sin K_{0}/8=-1$, $m=0.05$.
We further examine the effective monopole pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ in Fig. 10. Here, we have computed the band-projected pairing gap function in Eq. (16), using the tight-binding model given in Eq. (83) and projecting onto $\mathrm{FS}_{1}$ of the band Hamiltonian in Eq. (84). As depicted in Fig. 10 (a-b), the magnitude of the gap function reaches its maximum near the equator and vanishes at the poles, consistent with the effective pairing gap function in Eq. (37), calculated for the related continuum model. There are two gap nodes located at the north and south pole, which each display a local effective chiral-$f$ wave symmetry but with opposite chirality. Fig. 10 (a) and (c) are computed in the “north pole gauge” in which the Dirac string pierces the Fermi surface at $\theta_{\tilde{k}}=\pi$. In this gauge, the $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ shows a local $f+if$ pairing, as evident by the winding in Fig. 10 (c). In contrast, Fig. 10 (b) and (d) are computed in the “south pole gauge,” with the Dirac string piercing through $\theta_{\tilde{k}}=0$. In this gauge, the $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ is described by a local $f-if$ pairing, as shown in Fig. 10 (d). The $\mathrm{U}(1)$ topological obstruction due to the nonvanishing pair monopole charge $q_{p}=3$ leads to nonvanishing net vorticity, given by twice the pair monopole charge $62q_{p}=$ 6. The global angular momentum, which is a conserved quantity, is given by $l_{z,\mathrm{glob}}=0$. Moreover, the system with $\Delta_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}$ pairing order exhibits three chiral surface modes in the $ab$ plane under open boundary conditions. Although the local angular momentum $l_{z,\mathrm{loc}}$ of the pairing order changes as $k_{z}$ is varied between the gap nodes, as shown in the projection of the pairing gap function in Fig. 10, the leading term in the surface state dispersion is of order $\mathcal{O}(k^{2})$ and consequently overshadows terms of order $\mathcal{O}(k^{3})$ which encode the shift in local angular momentum. Nonetheless, the three surface modes are analogous to that which would arise in a system with chiral $f$-wave pairing.
Appendix F Monopole superconducting order with different $l_{z,\mathrm{glob}}$
Figure 11: (a) Josephson junction along the $z$ direction between a $\Delta_{\mathrm{MSC}}^{(q_{p}-1,l_{z}=2)}$ monopole superconductor and a $d_{x^{2}-y^{2}}+id_{xy}$-wave superconductor. (b) Josephson current phase relation between the $\Delta_{\mathrm{MSC}}^{(q_{p}-1,l_{z}=2)}$ monopole superconductor and $d_{x^{2}-y^{2}}+id_{xy}$-wave superconductor (red, solid), which shows nonvanishing first order coupling. In contrast, the first order Josephson current phase relation for a junction with an $s$-wave SC (black, dashed), is zero.
We demonstrate the nonvanishing first order Josephson critical current for a junction between a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=2)}$ monopole superconductor and a $d_{x^{2}-y^{2}}+id_{xy}$-wave superconductor, as detailed in Sec. IV.2. We consider a junction along the $z$ direction, shown schematically in Fig. 11 (a) and compute the Josephson current using a tight-binding model of the Josephson junction with spin-independent hopping at the junction link, analogous to that described in Sec. IV.1.2. The Josephson current phase relation is shown in red in Fig. 11 (b). For the same symmetry reasons as discussed in Sec. IV.2, there is nonvanishing first order Josephson current between a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=2)}$ monopole superconductor and $d_{x^{2}-y^{2}}+id_{xy}$ superconductor, i.e. both pairing transform under rotation $R_{z}$ according to global angular momentum $l_{z}=2$, there is nonvanishing first order Josephson coupling. The magnitude of the Josephson current is largely suppressed due to the gap nodes along the $k_{z}$ axis for both the chiral $d$-wave and monopole superconductor. Nonetheless, the $2\pi$-periodic first order Josephson current persists, indicative of the global angular momentum of the monopole superconducting pairing order, $l_{z}=2$. This is in contrast to, for example, a junction with a junction between an $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=2)}$ monopole superconductor and $s$-wave superconductor, for which the total Josephson current would cancel upon summing the contributions from states with different conserved transverse momentum $k_{\parallel}$.

### v2

@@ E.1 Tight-binding model of $q_{p}=-1$ monopole superconductor
Figure 10: The magnitude [(a) and (b)] and phase winding [(c) and (d)] of the monopole pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ based on the tight-binding model in Appendix E.1. The left and right columns respectively correspond to the north and south pole views of the pairing order at FS1. (c) and (d) are plotted under the gauge that the pairing phase is well-defined at the north and south pole respectively. Here, we use the same parameters as those for Fig. 9.
E.2 Tight-binding model of $q_{p}=3$ monopole superconductor
@@
We construct a tight-binding model for the monopole superconducting order having pair monopole charge $q_{p}=3$, as introduced in Sec. IV.3. Here, the band Hamiltonian describes a magnetic doped triple Weyl semimetal. As $C_{6}$ symmetry is needed to stabilize the triple Weyl nodes, the tight-binding model is defined on a three-dimensional hexagonal lattice with $D_{6h}$ symmetry, which consists of $AA$ stacked two-dimensional triangular lattices.
@@
$H_{\mathrm{MSC}}^{(q_{p}=3,l_{z}=0)}=\sum_{\mathbf{k}}\sum_{\sigma,\sigma^{\prime}=\uparrow,\downarrow}c^{\dagger}_{\mathbf{k},\sigma}\left[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}(\mathbf{k})\right]_{\sigma,\sigma^{\prime}}c_{\mathbf{k},\sigma^{\prime}}$
$+\sum_{\mathbf{k}}\sum_{\sigma,\sigma^{\prime}=\uparrow,\downarrow}\Big{(}c^{\dagger}_{\mathbf{k},\sigma}\big{[}\Delta_{\mathrm{inter}}(\mathbf{k})\big{]}_{\sigma,\sigma^{\prime}}c^{\dagger}_{-\mathbf{k},\sigma^{\prime}}+\mathrm{h.c.}\Big{)},$ (88)
with $\Delta_{\mathrm{inter}}(\mathbf{k})$ being the inter-Fermi surface pairing. Here, the Fourier transform of the band Hamiltonian kernel for the triple Weyl semimetal is
@@
where $g_{x}(\mathbf{k})=2t_{1}(\sin k_{1}-\sin k_{2}+\sin k_{3})$, $g_{y}(\mathbf{k})=({-2t_{2}}/{3\sqrt{3}})[\sin(k_{1}+k_{2})-\sin(k_{2}+k_{3})+\sin(k_{3}-k_{1})]$, and $g_{z}(\mathbf{k})={2t_{z}(\cos K_{0,z}-\cos k_{z})}+{2m(3-\cos k_{1}-\cos k_{2}-\cos k_{3})}$. The above system hosts two triple Weyl nodes pinned along the $k_{z}$ axis at ±$\pm\mathbf{K}_{0}=\left(0,0,\pm K_{0,z}\right)$ with opposite winding numbers $\mp 3$. In the low-energy regime, the above tight-binding model of the band Hamiltonian reduces to that of the continuum model given in Eq. (36). Nearest and next-nearest neighbor hopping amplitudes are denoted by $t_{1}$ and $t_{2}$ respectively which, for simplicity, we choose to be $t_{1}=t_{2}=8t_{z}\sin K_{0,z}$ to compare to the continuum model. Here, $k_{i}\equiv\bm{\delta}_{i}\cdot\mathbf{k}$, where $\bm{\delta}_{1}=(1,0,0)$, $\bm{\delta}_{2}=(\frac{1}{2},\frac{\sqrt{3}}{2},0)$, and $\bm{\delta}_{3}=(-\frac{1}{2},\frac{\sqrt{3}}{2},0)$ are the lattice vectors connecting neighboring sites in the $ab$-plane, and the lattice constant has been set to unity. At appropriate doping, the above system realizes two disjoint Fermi surfaces, $\mathrm{FS}_{1}$ and $\mathrm{FS}_{2}$, situated about triple Weyl nodes at ±$\pm\mathbf{K}_{0}$ and related by parity, $\mathcal{H}_{\mathrm{kin},1}^{(\nu=-3)}(\mathbf{k})=\sigma_{z}\mathcal{H}_{\mathrm{kin},2}^{(\nu=+3)}(-\mathbf{k})\sigma_{z}$.
@@
The nonzero matrix elements, which feature nearest and next-nearest neighbor hopping, read
n;n $=(2t_{z}\cos K_{0,z}+6m)\sigma_{z}-\mu\sigma_{0},$ (90)
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{1}}$ $=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{1};\mathbf{n}}^{\dagger}=-m\sigma_{z}-it_{1}\sigma_{x}$
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{2}}$ $=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{2};\mathbf{n}}^{\dagger}=-m\sigma_{z}+it_{1}\sigma_{x}$
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{3}}$ $=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{3};\mathbf{n}}^{\dagger}=-m\sigma_{z}-it_{1}\sigma_{x}$
@@
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{1}+\bm{\delta}_{2}}$ $=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{1}+\bm{\delta}_{2};\mathbf{n}}^{\dagger}=-\frac{it_{2}\nu}{3\sqrt{3}}\sigma_{y}$
@@
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{2}+\bm{\delta}_{3}}$ $=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{2}+\bm{\delta}_{3};\mathbf{n}}^{\dagger}=\frac{it_{2}\nu}{3\sqrt{3}}\sigma_{y}$
@@
$[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n};\mathbf{n}+\bm{\delta}_{3}-\bm{\delta}_{1}}$ $=[\mathcal{H}_{\mathrm{kin}}^{\mathrm{TWSM}}]_{\mathbf{n}+\bm{\delta}_{3}-\bm{\delta}_{1};\mathbf{n}}^{\dagger}=-\frac{it_{2}\nu}{3\sqrt{3}}\sigma_{y}$
[…]